OpenStudy (anonymous):

Verify the equation below with each of the values listed for x to find a solution. 5 = 3x + 4

OpenStudy (anonymous):

Help

OpenStudy (anonymous):

a) subtract 4 b) divide by 3

OpenStudy (anonymous):

2 = 3x + 4; -3x = 4 - 2; -3x = 2; x = -2/3; 2 = 3(-2/3) + 4; 2 = -2 + 4; 2 = 2